A 7% semi-annual coupon bond settles 12/11/15 at 105.00 (clean price). Bond maturity is 12/31/25. For a 300 basis points increase in yield, assuming a 30/360-day count convention, what is the actual (dirty) price?

  • 86.72
  • 88.19
  • 83.61
  • 82.53
  • 85.60
Answer #1

Actual price = 88.19

Step 1). Calculate the current YTM using the given clean price of 105.

Step 2). New YTM = current YTM + 3%

Step 3). Calculate clean price, as on the settlement date and accrued interest from last coupon payment date to settlement date.

Step 4). Add clean price and accrued interest to get the actual price.

30-06-2015 11-12-2015 31-12-2025 7% 151 Last coupon payment date 152 Settlement date 153 Maturity date 154 Coupon rate 155 Cl

Formulas:

151 Last coupon payment date =DATE(2015,6,30) 152 Settlement date =DATE(2015,12,11) 153 Maturity date =DATE(2025,12,31) 154 C

Note: Keep in mind that PRICE() function calculates the clean price, not the dirty price.